32 Portfolio Management A portfolio manager is concerned that two stocks generate an annual income of $15,000 for a client. The two stocks earn annual dividends of $2.40 and $3.50 per share, respectively. If x equals the number of shares of stock 1 and y equals the number of shares of stock 2, the equation
2.4x+3.5 15,000
states that total annual dividend income from the two stocks should total $15.000 Rewrite the equation in slope-intercept form and interpret the meaning of the slope and y intercept in this application. Solve for the x intercept and interpret its meaning.

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

the equation is 2.4x + 3.5y = 15000
x = number of shares of stock 1.
y = number of shares of stock 2.
this is a straight line equation in standard form.
convert to slope intercept form by solving for y.
subtract 2.4x from both sides of the equation to get 3.5y = 15000 - 2.4x
divide both sides by 3.5 to get y = 15000/3.5 - 2.4x/3.5
rearrange the equation in descending order of degree to get y = -2.4x/3.5 + 15000/3.5
this is the same as y = -2.4/3.5 * x + 15000/3.5
the equation is now in slope intercept form of y = mx + b
m is the slope
b is the y-intercept.
m = -2.4/3.5
b = 15000/3.5
the y-intercept is the value of y when x = 0.
y = -2.4/3.5 * 0 + 15000/3.5 which results in y = 15000/3.5 = 4285.714286 which is equal to 4285.714 rounded to 3 decimal places.
the x-intercept is the value of x when y = 0 which results in 0 = -2.4/3.5 * x + 15000/3.5
subtract 15000/3.5 from both sides of the equation to get -15000/3.5 = -2.4/3.5 * x
solve for x to get x = (-15000/3.5) / (-2.4/3.5) = 6250.
the equation can be graphed as shown below.

Answer by ikleyn(52781)

.

Looking at this post,  I see that the basic equation is written incorrectly in it
and the assignments are quite non-sensical.


It is clearly seen from the post, that the person who composed it,
has no right idea on how to make a normal Math problem from it,
as well as,  has no right idea what to do with these words, that jump out of his mouth, at all.